OpenStudy (anonymous):

Need help with a problem from my homework. Express answer in exact form. Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. (Hint: A chord divides a circle into two segments. In problem 1, you found the area of the smaller segment.)

OpenStudy (anonymous):

The answer should be like this

OpenStudy (anonymous):

Find the area of of the larger segment by finding the area of the whole circle and subtracting the answer from question 1. Area of the circle = pi*r^2 Area of the circle = 3.14*8^2 = 301.06 Area of the larger segment = 201.06 - 5.798 = 195.26

OpenStudy (anonymous):

The answer needs to actually be like that equation \[A=(\frac{ [][][] }{ [] }\pi + [][]\sqrt{[]})\]

OpenStudy (anonymous):

You just plug that information in...

OpenStudy (anonymous):

Im bad at what I do ok? their is a reason I am in summerschool right now

OpenStudy (anonymous):

Sorry sorry lemme see...

OpenStudy (anonymous):

We know that the triangle formed by the two radii and the chord is an equilateral triangle, all angles are 60 degrees, which is 1/6 of 360 degrees Find area inside the 60 degree arc 1%2F6*pi%2A8%5E2 = 35.51 Find the area of the equilateral triangle 1%2F2*8*sqrt%288%5E2-4%5E2%29 = 27.71 sq/in Find the area of the shape enclosed by the 60 degree arc and the chord 35.51 - 27.71 = 7.8 sq/in Find the area of the larger segment pi%2A8%5E2 - 7.8 = 193.26 sq/inches

OpenStudy (anonymous):

pi*8^2- 7.8 = 193.26 sq/inches

OpenStudy (anonymous):

^^^^^^^short to the point answer